1)A canoeist paddles directly across a river at 5 m/s. At the same time, the river is pushing the canoeist sideways at 3 m/s. What is the angle of the canoe's motion relative to the riverbank?

2)A molecule in the air has a mass of 2 x 10-25 kg. At a temperature of 25ºC, how fast is it moving?
Answer #1

1(A)

if we take river velocity in positive x direction and canoeist velocity in y direction . then

angle = tan-(5/3) = 59.036 from positive x direction

2(A) it will V_rms = \sqrt{3RT/M}

6.927*1014